OpenStudy (anonymous):

HELP!! 2x^3(5x^3 - 7)

OpenStudy (debbieg):

Start by distributing. \(a(b-c)=ab-ac\) So apply that with \(a=2x^3, b=5x^3 \text{ and } c= 7\)